OpenStudy (anonymous):

Find magnitude and direction of the vector I will medal and fan please thanks!

hartnn (hartnn):

can you post the pic here? it is too small to read.. Although, you can start: we have x and y component. \(magnitude = \sqrt{x^2+y^2} \\ direction = \angle \theta = tan^{-1}(\dfrac{y}{x})\)

hartnn (hartnn):

ok, thats fine. x = 75 y = 180 use the formula and try to get the magnitude and angle. let me know if you get stuck.

hartnn (hartnn):

better :)

hartnn (hartnn):

good!

hartnn (hartnn):

\(\tan^{-1}(y/x) = \tan^{-1}(180/75)\) which, when you use calculator, you get 67 degrees approx.

hartnn (hartnn):

and the direction is between south and east. 195mi, 67 south of east :)

OpenStudy (anonymous):

Ohh I now get it thank you

OpenStudy (anonymous):

Find cos t for this triangle wait one sec

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

ok

hartnn (hartnn):

which angle is t?

hartnn (hartnn):

cos t = side adjacent to t/hypotenuse

OpenStudy (anonymous):

Yea I post pic

OpenStudy (anonymous):

This one is confusing with the square roots

hartnn (hartnn):

you can post the pic here, using "Attach File" button

hartnn (hartnn):

adjacent side is 8, hypotenuse = 16 8/16 = ..?

OpenStudy (anonymous):

I don't have that button 8/16 is cos T but answer choices are 1/2, 1/sqrt3, sqrr3/1, and swrt3/2

hartnn (hartnn):

16 = 2*8

hartnn (hartnn):

and everyone has the "Attach File" button. Its on the left of "Post" button below

OpenStudy (anonymous):

I'm using mobile

hartnn (hartnn):

okk.. 8/16 = 8/(2*8) 8 cancels out. 1/2

OpenStudy (anonymous):

Oh 1/2 since 8 cancels out, thanks! And I have some couple of questions more

hartnn (hartnn):

post them as a new question, so that if i am unavailable, others can answer them :)

OpenStudy (anonymous):

Determine if triangles these are similar and if they are what postulate or theorem proves true similarity, profile pic is the diagram, thanks and ok:))